What volume of O2(g) measured at 1 atm and 273 K will be formed by action of 100 ml of 0.5 N KMnO4 on hydrogen peroxide in an acid solution? The skeleton equation for the reaction is 
KMnO4 + H2SO4 + H2O2 → K2SO4 + MnSO4 + O2 + H2O

Grade:12th pass

1 Answers

User
5 years ago
Equating gram-equivalent of KMnO4 and O20.1×0.5= vol.of O2/5.6Vol.of O2 =0.05×5.6 = 0.28 litreMind it that gram equivalent of gases =vol. Of gas at STP /equivalent volume at STP.Since, 32 gm O2=22.4 litre So,8gm O2=5.6 litre